OpenStudy (anonymous):

Wright an equasion of a line that passes through the point (-2,8) and is PERPENDICULAR to the line represented by this 5x-2y=-10

OpenStudy (anonymous):

1º write 5x-2y=-10 in slope y-intercept form 2º take the negative inverse of the slope , call it S 3º put this slope S from 2º in this equation (y-P)=S(x-Q), where(P,Q) is the point you whant you line to go through

OpenStudy (anonymous):

May i have the answer lol srry if im bugging yu

OpenStudy (anonymous):

dude, make some efort

OpenStudy (anonymous):

Jsut this one :P

OpenStudy (anonymous):

the previous was exactly the same

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

PLZ

OpenStudy (anonymous):

just try it, it's to easy ...

OpenStudy (anonymous):

go step by step

OpenStudy (anonymous):

is it 2x-5y=26 or -2x+5y=-13

OpenStudy (anonymous):

do 1º step

OpenStudy (anonymous):

ahh i figued either this or that one 2x-5y=26 or -2x+5y=-13 Come on!

OpenStudy (anonymous):

sry, can't help you like that

OpenStudy (anonymous):

i got the frst one

OpenStudy (anonymous):

?

OpenStudy (anonymous):

show

OpenStudy (anonymous):

ahh its alot

OpenStudy (anonymous):

Just a simple yes or no

OpenStudy (anonymous):

plz